A proton moves in a circle of radius 1.26m in a re
Answer #1


10)


IN the magetic field


Fc = Fb


mv^2/r = qvB

r = mv/qB


timeperiod T = 2pir/v = 2pim/qB


frequency f = 1/T = qB/(2*pi*m) =


f = (1.6*10^-19*90*10^-3)/(2*pi*1.67*10^-27)

f = 1.37 MHz   <<<<<<<---------answer

+++++++++++++++++


kinetic energy

K = 0.5*m*v^2

r = mv/qB


v = rqB/m

k = 0.5*r^2*q^2*B^2/m

k = 0.5*1.26^2*(1.6*10^-19)^2*0.9^2/(1.67*10^-27)


k = 9.86*10^-12 J

k = 9.86*10^-12/(1.6*10^-19)


k = 61.6 MeV